Use the properties of logarithms to simplify each equation and solve it.
Round to the nearest hundredth.


49. ln (x - 3) = 1
50. 2 ln x - ln 3 = 2
51. ln (x + 2) + ln (x - 2) = 0

Answer by Alan3354(21583)

Use the properties of logarithms to simplify each equation and solve it.
Round to the nearest hundredth.

49. ln (x - 3) = 1
x - 3 = e
x = e+3
x = ~ 5.72
-----------------
50. 2 ln x - ln 3 = 2
ln(x^2/3) = 2
x^2/3 = e^2
x^2 = e^2/3
x = +e*sqrt(3)/3
x = -e*sqrt(3)/3
------------------
51. ln (x + 2) + ln (x - 2) = 0
ln((x+2)/(x-2)) = 0
(x+2)/(x-2) = 1
x+2 = x-2 --> inconsistent, no solution.